# On the use of Graphs for Satellite Image Time Series

### v1

@@ 3.3 Fidelity and complexity
Despite the advantages of a graph-based approach, modeling SITS as an intermediate step necessarily induces additional complexity for data representation compared to direct data use. However, the processing of this representation can be less costly, leading to a compromise between the accuracy of the representation and its processing. It is therefore important to know the cost of such a representation in terms of model fidelity, storage and computational complexity.
@@ 3.3.1 Model fidelity
An ideal embedding from a fidelity point of view should contain the same information as in its corresponding object, both radiometric, textural and geometrical ones. However, it is also crucial for an embedder to extract higher-order features that are useful for the downstream task. A good embedding must be a compromise between an embedding capable of reconstructing the region identically and a lossy embedding with generalization capabilities. For example, a mean aggregator is not a high-fidelity encoder as it loses a lot of information, but focuses on some part of the radiometric data that could be crucial for specific applications. Recently, learning methods for encoding objects have given good results (Section 3.2.1), but the level of fidelity of embeddings and their ability to generalize is more difficult to assess than for deterministic methods, as they depend on the training data.
@@ 4.4 Intrinsic regression using graph
To illustrate the benefits of graphs for these tasks, the works cited in this section cover those using SITS-derived data. These may include spatio-temporal reanalysis of the atmospheric state, landslide displacement based on sequences of point clouds obtained by radar interferometry, or wildfire evolution maps. Those SITS-derived data face the same challenges than SITS regarding the processing of spatial and temporal dimensions [52]. Thus, in order to illustrate the value of graphs for intrinsic regression, despite the paucity of work with raw SITS, we are relying in addition on these SITS-derived data.
@@ 5.1.1 Data
DynamicEarthNet consists of a dataset of 75 SITS taken from all around the world. Figure 11 shows the location of each area, along with a few example SITS. The SITS acquired from January 2018 to December 2019 are composed of multispectral (RGB + near-infrared reflectance) images from the Planet Fusion product of size 1024 $\times$ 1024 pixels at high spatial (3 meters per pixel) and temporal (daily) resolutions. A land cover ground truth is provided at the beginning of each month (i.e., 24 labeled images among the 730 per SITS) by manual annotation among 7 classes: 7.1% impervious surfaces, 10.3% agriculture, 44.9% forest & other vegetation, 0.7% wetlands, 28.0% soil, 8.0% water, 1.0% snow. To keep this use case easily reproducible, we choose to restrict the data to annotated images only, diving our task to the fully supervised setting. It is still possible to extend the method to a semi-supervised setting by keeping the unlabeled daily acquisitions, thus enhancing the temporal consistency.
@@ 5.1.3 Experimental setup
To study the contribution of spatio-temporal graphs in the analysis of SITS, we compare GNNs to a pixel-based model as baseline, U-Net [221]. It exploits only the spatial context at multiple scales and produces a segmentation map of results independently for each date. The graph method is evaluated with different graph convolutional layers to highlight the importance of GNN design choices related to the task, namely GCN [144], GraphSAGE [222], GATv2 [223] and ResGatedGCN [224]. We also include a multi-layer perceptron (MLP), which acts as the previous GNNs but on edge-free graphs, i.e., without consideration of the spatio-temporal context, following a traditional OBIA framework.
GNN architectures consist of an encoder of four graph convolution layers, with an hidden dimension of 64, followed by a linear layer for classification. To aggregate information from both the spatial and temporal dimensions, the encoder consists of two layers aggregating only from the spatial neighbors, and two layers using temporal neighbors, as in Figure 9. Each convolution layer is followed by a batch normalization and a ReLU activation. U-Net experiments are run under the same setting as in the DynamicEarthNet paper [220]. Thus, we refer to this work for the configuration of pixel-based baseline and to our implementation for other details about the setup. Another U-Net setting, denoted U-Net/SITS, is also evaluated for which an input batch contains images from only the same SITS, rather than single images extracted from multiple SITS. This setting provides a similar data loading scenario than the one used with GNNs. Although it is not optimal for U-Net by mitigating the learning of varied data in the same batch, this setting helps us to compare spatio-temporal and spatial-only methods more fairly by circumventing the limitation of input data size.
The Felzenszwalb algorithm [81] used to obtain objects relies on the Scikit-Image [225] implementation with $scale=100$ and $min\_size=5$ hyperparameters. All the deep learning models are trained using Adam optimizer on cross-entropy loss with a learning rate of 10-4. The training process, for each model, is conducted over 100 epochs with training batch size equal to 1 for GNNs, 4 for U-Net and 24 for U-Net/SITS. Detailed hyperparameter values are available for all models at github.com/corentin-dfg/graph4sits/.
@@
Training and testing are carried out on Intel® Cascade Lake 6226 12 cores@2.7GHz with 384 GB of RAM and an Octo-GPU v100. Preprocessing is done on CPU 12th Gen Intel® Core™ i7-12800H 20 cores@2.4-4.8GHz. Networks are implemented using PyTorch and PyG.
@@ 5.1.4 Experimental results
Table 1 summarizes the average quantitative performances on DynamicEarthNet achieved by the pixel-based baseline U-Net, MLP and GNNs. In the following, we first highlight the role of context information for the case study compared to a context-agnostic model, such as MLP. Then, we inspect the GNN performances related to the pixel-based baseline. We also analyze the inherent error of object-based methods that results from the semantic segmentation step. Finally, we inspect the computational performance of the models in terms of training and testing time.
@@
The GraphSAGE model outperforms MLP on accuracy and IoU metrics, demonstrating the value of using the spatio-temporal context of an object in addition to its own features to determine its nature. These performances are significantly better for the classes agriculture, water, and especially impervious surface. These are the classes for which determination by spectral characteristics is limited, either because of the very nature of these classes (e.g., agriculture is confused with forest and soil) or because of high intra-class variability. Indeed, impervious surface is confused with forest, soil and water. A possible source of this confusion is because geographical objects in the impervious surface class are generally smaller and more difficult to segment into regions. The gap between MLP and GraphSAGE performance thus shows that learning the context of each object decreases the class ambiguity.
@@
Improvements over convolutional GNNs include attention (GATv2) or gate (ResGatedGCN) mechanisms, which enable anisotropic aggregation of neighbor information. From a message-passing point of view (see Section 4.3.2), self-attention consists in weighting incoming messages according to the similarity of the target node with the source nodes. The gate mechanism is an independent weighting of each incoming message feature, still depending on the characteristics of the source and target nodes. Contrary to expectations, the performance of these two networks is inferior to that of a GNN with isotropic aggregation such as GraphSAGE, which treats all its neighbors in the same way. Regarding GATv2, even though the benefits of dynamic context awareness are found for impervious surface and especially water, it seems that performances for forest & other vegetation and soil are lower than those of GraphSAGE. Despite visual analysis and post-investigations, we were not able to explain this unexpected behavior, which thus requires further analysis. With ResGatedGCN, there is a drop in performance compared to GraphSAGE and even to MLP. This does not seem to be due to anisotropic aggregation—otherwise a similar observation would be made with GATv2—but rather to the way edge gates work. This doubles the number of parameters required and makes convergence more difficult when learning gates, despite their greater expressiveness (in fact, gates can be seen as a soft attention process). The less frequent the class, the lower the performance, suggesting that this model requires more data to be correctly trained.
This analysis helps explain why GraphSAGE is a better choice of graph convolution for the studied task and data.
@@
Data batch processing is another criterion influencing the performance of U-Net and GNNs. Indeed, as U-Net uses only the spatial dimension and works by date, it is easier to integrate different locations in the same batch. This makes it easier to take into account data variety and class distribution at each network iteration. On the other hand, the use of spatio-temporal graphs restricts the size of the input batch due to the volume of each graph, recalling that each graph in the current experiment is built from 24 images. The input data at each iteration is therefore less representative of the dataset as a whole, resulting in slower and less stable convergence. To illustrate this phenomenon, the U-Net model was trained with a batch size of 24 corresponding to one SITS per iteration, thus approximating the data loading used with GNNs. The results are reported in Table 1 as U-Net/SITS and clearly demonstrate the importance of data representativeness in a batch. The trade-off between representativeness and input data size is crucial in spatio-temporal models. Comparing the results of the U-Net/SITS model trained on a similar setting than the graph-based models shows the value of a spatio-temporal approach.
@@ 5.1.5 Discussion
This case study highlights a practical downstream task by illustrating the different design and processing questions which can emerge by using spatio-temporal graphs for SITS semantic segmentation over time. The main purpose of spatio-temporal graphs lies in the integration of spatial and temporal information into the same data structure, thus enabling the direct use of spatio-temporal context. They also limit the data redundancy, as the processing is based on homogeneous regions, thus reducing computational costs. However, despite these qualities, graph methods are understudied, and the design choices are poorly explained even for specific applications. To enable better use of graph structure for SITS, we encourage the scientific community to further study these issues to achieve an efficient pipeline for classification tasks, particularly with a view to matching the performance of pixel-based methods. A number of avenues have been highlighted in the previous sections, including optimization of object segmentation (e.g., learned segmentation, 3D spatio-temporal segmentation of the datacube), improvement of deep feature extraction for irregular objects, and study of the properties of the SITS-based graphs (e.g., heterophily [226], connectivity, heterogeneity, class balancing) to subsequently guide design and processing choices. Concerning the learning part, transferring the latest advances from graph deep learning is necessary, both in terms of learning strategy (subgraph batch training [222, 227]) and architecture (deeper GNN [228], graph transformers [229], space-first time-later vs. time-first space-later vs. joint space-time [230, 231]). Finally, it is important that design and architecture choices in future work are explicitly motivated, and the use of interpretability tools looks promising for these purposes. The current maturation phase of spatio-temporal graphs applied to SITS is crucial, and we hope that this case study may aid future work in classification and semantic segmentation.
@@ 5.2.4 Experimental results
Input average - $0.1550$ $23.32$ $0.7465$
Persistence - $0.1332$ $25.03$ $0.7897$
@@
Figure 15: RMSE bar plot depending on the targeted residuals since the last acquisition for each pixel of the test set. The background color represents the training pixel distribution of the targeted residuals. Targeted residuals can theoretically range from $-2$ to $2$, but this does not happen in practice in the dataset.
@@
Last-date SLIC 0.1097$\pm$0.0035 26.42$\pm$0.27 0.8170$\pm$0.0070
Multi-date SLIC 0.1101$\pm$0.0034 26.12$\pm$0.45 0.8151$\pm$0.0076
@@ 5.2.5 Discussion
This case study illustrates another downstream task that still uses graphs, albeit with different role, design and processing choices. SITS forecasting is a pixel-to-pixel task, thus, in contrast to the previous case study, graphs are used both to link pixel and object scales and to model inter-object relationships. Indeed, it brings a prior spatial information using geographical objects to learn local dynamics of water basins and their surroundings. In this work, we use a simple region adjacency graph, which prevents to capture long-range dependencies. To improve generalization capability of the model in larger images, the proposed architecture should be extended to handle multiple scales, e.g., using multi-meshes or hierarchical meshes. Besides, the segmentation step giving prior spatial information to the model still results from an arbitrary setting, i.e., the number of superpixels. Other segmentation algorithms, which can produce different types of regions, could influence performance, and thus optimizing this object segmentation step is a means of improvement as other hyperparameters of graph design. Moreover, temporal information is not processed using a graph, but embedded at pixel level with a simple MLP not tailored to learn temporal patterns from time series. This results from the need to use a static spatial graph to represent the prior spatial information which is considered constant. It appears as a current limitation for highly dynamics phenomena where relationships between the object can vary. Solutions including time-adaptive meshes [238, 198, 239] and discrete-time dynamic graphs [240] need to be explored in order to retain the representativeness of graphs to model a time-evolving scene. Finally, as in the previous study, interpretability of models and results is crucial to motivate further developments. A way to improve it is realizing a more robust analysis of the results by considering disaggregated results as done in Figure 14 and Figure 15. The distinction can be done on various criteria which add semantics, such as seasonality and rareness of the events, classification of the phenomenon or of the geographical site. Therefore, such a benchmark should enhance the evaluation and comparison of different models and help explain the pros and cons of each model. We hope that this case study may help future works to better consider advantages and current limitations of graphs for a SITS forecasting task, thus encouraging the development of new methods.
@@ 6 Challenges and Perspectives
The abundance of data in Earth observation is the main cause of the scalability problem, with a constant flow of new data which are sparsely labelled. Various propositions of solution to handle this data volume exist, such as foundation models [241, 242] and continual learning. Exploring these solutions in the setting of spatio-temporal graphs could be promising.
@@
In this paper, graphs are principally used as tool to analyze and process SITS from a single source, mainly optical. However, Section 3.1 gave an overview of the wide range of sensors imaging the Earth. This diversity of sources brings development opportunities to adapt to each setting, even those with little data currently available (hyperspectral SITS, off-nadir and multi-angle acquisitions, 3D point clouds sequence). Graph-based approaches can also be adapted for data similar to SITS, such as aerial or drone imagery and numerical weather simulations.

### v2

@@ 3.3 Fidelity and complexity
Despite the advantages of a graph-based approach, modeling SITS as an intermediate step necessarily induces additional complexity for data representation compared to direct data use. However, the processing of this representation can be less costly, leading to a compromise between the accuracy of the representation and its processing. It is therefore important to know the cost of such a representation in terms of model fidelity, storage and computational complexity. The theoretical complexities reported in this section are consistent with experiments, both in terms of runtime and memory usage. For space reasons, these experiments are not reported in the paper, but are available on request from the corresponding author.
@@ 3.3.1 Model fidelity
An ideal embedding from a fidelity point of view should contain the same information—radiometric, textural, and geometric—as in its corresponding object. However, it is also crucial for an embedder to extract higher-order features that are useful for the downstream task. A good embedding must be a compromise between an embedding capable of reconstructing the region identically and a lossy embedding with generalization capabilities. For example, a mean aggregator is not a high-fidelity encoder as it loses a lot of information, but focuses on some part of the radiometric data that could be crucial for specific applications. Recently, learning methods for encoding objects have given good results (Section 3.2.1), but the level of fidelity of embeddings and their ability to generalize is more difficult to assess than for deterministic methods, as they depend on the training data.
@@ 4.4 Intrinsic regression using graph
To illustrate the benefits of graphs for these tasks, the works cited in this section cover those using SITS-derived data. These data may include spatio-temporal reanalysis of the atmospheric state, landslide displacement based on sequences of point clouds obtained by radar interferometry, or wildfire evolution maps. Those SITS-derived data face the same challenges than SITS regarding the processing of spatial and temporal dimensions [224]. Thus, in order to illustrate the value of graphs for intrinsic regression, despite the paucity of work with raw SITS, we are relying in addition on these SITS-derived data.
@@ 5.1.1 Data
DynamicEarthNet consists of a dataset of 75 SITS taken from all around the world. Figure 11 shows the location of each area, along with a few example SITS. The SITS acquired from January 2018 to December 2019 are composed of multispectral (RGB + near-infrared reflectance) images from the Planet Fusion product of size 1024 $\times$ 1024 pixels at high spatial (3 meters per pixel) and temporal (daily) resolutions. A land cover ground truth is provided at the beginning of each month (i.e., 24 labeled images among the 730 per SITS) by manual annotation among 7 classes: 7.1% impervious surfaces, 10.3% agriculture, 44.9% forest & other vegetation, 0.7% wetlands, 28.0% soil, 8.0% water, 1.0% snow. To keep this use case easily reproducible, we choose to restrict the data to annotated images only, concentrating our task to the fully supervised setting. It is still possible to extend the method to a semi-supervised setting by keeping the unlabeled daily acquisitions, thus enhancing the temporal consistency.
@@ 5.1.3 Experimental setup
To study the contribution of spatio-temporal graphs in the analysis of SITS, we compare GNNs to a pixel-based model as baseline, U-Net [175]. It exploits only the spatial context at multiple scales and produces a segmentation map of results independently for each date. The graph method is evaluated with different graph convolutional layers to highlight the importance of GNN design choices related to the task, namely GCN [105], GraphSAGE [74], GIN [214], GATv2 [20] and ResGatedGCN [19]. We also include a multi-layer perceptron (MLP), which acts as the previous GNNs but on edge-free graphs, i.e., without consideration of the spatio-temporal context, following a traditional OBIA framework.
[…]
@@
Training and testing are carried out on Intel® Cascade Lake 6226 12 cores@2.7GHz with 384 GB of RAM and an Octo-GPU v100. Preprocessing is done on CPU 12th Gen Intel® Core™ i7-12800H 20 cores@2.4-4.8GHz. Networks are implemented as defined in their original their paper using PyTorch and PyG.
@@ 5.1.4 Experimental results
Table 1 summarizes the average quantitative performances on DynamicEarthNet achieved by the pixel-based baseline U-Net, MLP and GNNs. To highlight the limitations of the different models, we provide in Figure 12 complementary information with the confusion matrices over the validation set. In addition, part of the inter-class error is due to class imbalance in the dataset, which is suggested by the similar structure of each confusion matrix, regardless of the model. In the following, we first highlight the role of context information for the case study compared to a context-agnostic model, such as MLP. Then, we inspect the GNN performances related to the pixel-based baseline. We also analyze the inherent error of object-based methods that results from the semantic segmentation step. Finally, we inspect the computational performance of the models in terms of training and testing time.
@@
The GraphSAGE model outperforms MLP on accuracy and IoU metrics, demonstrating the value of using the spatio-temporal context of an object in addition to its own features to determine its nature. These performances are significantly better for the classes agriculture, water, and especially impervious surface. These are the classes for which determination by spectral characteristics is limited, either because of the very nature of these classes (e.g., agriculture is confused with forest and soil, as shown by Figure 12) or because of high intra-class variability. Indeed, impervious surface is confused with forest, soil and water. A possible source of this confusion is because geographical objects in the impervious surface class are generally smaller and more difficult to segment into regions. The gap between MLP and GraphSAGE performance thus shows that learning the context of each object decreases the class ambiguity.
@@
Some of the failures are also due to the neighborhood aggregator used. Indeed, the aggregation function is one of the most critical components of GNN as it designs which property of the neighborhood is determinant. For instance, neighborhood distribution is better captured by the mean aggregator, structure information, such as node degree, by the sum aggregator, and specific features of some neighbors by the max aggregator [214]. The difference between GraphSAGE and GIN, which are inherently similar except for their aggregation function (mean and sum, respectively), suggests that the neighborhood distribution is more important than the graph structure to classify regions. Therefore, GNNs that use a (weighted) mean aggregator, such as GCN, GraphSAGE and GATv2, perform better.
Improvements over convolutional GNNs include attention (GATv2) or gate (ResGatedGCN) mechanisms, which enable anisotropic aggregation of neighbor information. From a message-passing point of view (see Section 4.3.2), self-attention consists in weighting incoming messages according to the similarity of the target node with the source nodes. The gate mechanism is an independent weighting of the features of each incoming message, still depending on the characteristics of the source and target nodes. Contrary to expectations, the performance of these two networks is inferior to that of a GNN with isotropic aggregation such as GraphSAGE, which treats all its neighbors in the same way. Regarding GATv2, even though the benefits of dynamic context awareness are found for impervious surface and especially water compared to MLP, it seems that performances for forest & other vegetation and soil are lower than those of MLP and GraphSAGE. With ResGatedGCN, there is a drop in performance compared to GraphSAGE and even to MLP, similar to the GIN performance. This does not seem to be due to anisotropic aggregation—otherwise a similar observation would be made with GATv2—but rather to the actual aggregation function used. The weighting from edge gates is not normalized, unlike in GATv2, and the aggregator corresponds more to a weighted sum than a weighted average. As a result, despite the greater expressiveness of ResGatedGCN (in fact, gates can be seen as a soft attention process), GATv2 achieves better performance.
This analysis helps explain why GraphSAGE is a better choice of graph convolution for the studied task and data. This comparison allows us to determine which characteristics of graph convolution are important for the node classification of a spatio-temporal graph in our case study. It may help to build a more effective model in the future.
Figure 12: Confusion matrices on the validation set. The results are normalized over the targets and provided with average on three random initializations.
@@
[…]
@@ 5.1.5 Discussion
This case study highlights a practical downstream task by illustrating the different design and processing questions which can emerge by using spatio-temporal graphs for SITS semantic segmentation over time. The main purpose of spatio-temporal graphs lies in the integration of spatial and temporal information into the same data structure, thus enabling the direct use of spatio-temporal context. They also limit the data redundancy, as the processing is based on homogeneous regions, thus reducing computational costs. However, despite these qualities, graph methods are understudied, and the design choices are poorly explained even for specific applications. To enable better use of graph structure for SITS, we encourage the scientific community to further study these issues to achieve an efficient pipeline for classification tasks, particularly with a view to matching the performance of pixel-based methods. A number of avenues have been highlighted in the previous sections, including optimization of object segmentation (e.g., learned segmentation, 3D spatio-temporal segmentation of the datacube), improvement of deep feature extraction for irregular objects, and study of the properties of the SITS-based graphs (e.g., heterophily [127], connectivity, heterogeneity, class balancing) to subsequently guide design and processing choices. Concerning the learning part, transferring the latest advances from graph deep learning is necessary, both in terms of learning strategy (subgraph batch training [74, 33]) and architecture (aggregation functions [41, 23], deeper GNN [117], graph transformers [50], space-first time-later vs. time-first space-later vs. joint space-time [25, 58]). Finally, it is important that design and architecture choices in future work are explicitly motivated, and the use of interpretability tools looks promising for these purposes. The current maturation phase of spatio-temporal graphs applied to SITS is crucial, and we hope that this case study may aid future work in classification and semantic segmentation.
@@ 5.2.4 Experimental results
Input average - $0.1550$ $23.32$ $0.7465$ -
Persistence - $0.1332$ $25.03$ $0.7897$ -
@@
Figure 16: RMSE bar plot depending on the targeted residuals since the last acquisition for each pixel of the test set. The background color represents the training pixel distribution of the targeted residuals. Targeted residuals can theoretically range from 2-2 to 22, but this does not happen in practice in the dataset.
@@
[…]
@@ 5.2.5 Discussion
This case study illustrates another downstream task that still uses graphs, albeit with different role, design and processing choices. SITS forecasting is a pixel-to-pixel task, thus, in contrast to the previous case study, graphs are used both to link pixel and object scales and to model inter-object relationships. Indeed, it brings a prior spatial information using geographical objects to learn local dynamics of water basins and their surroundings. In this work, we use a simple region adjacency graph, which prevents to capture long-range dependencies. To improve generalization capability of the model in larger images, the proposed architecture should be extended to handle multiple scales, e.g., using multi-meshes or hierarchical meshes. Besides, the segmentation step giving prior spatial information to the model still results from an arbitrary setting, i.e., the number of superpixels. Other segmentation algorithms, which can produce different types of regions, could influence performance, and thus optimizing this object segmentation step is a means of improvement. Moreover, temporal information is not processed using a graph, but embedded at pixel level with a simple MLP not tailored to learn temporal patterns from time series. This results from the need to use a static spatial graph to represent the prior spatial information which is considered constant. It appears as a current limitation for highly dynamics phenomena where relationships between the object can vary. Solutions including time-adaptive meshes [8, 162, 85] and discrete-time dynamic graphs [219] need to be explored in order to retain the representativeness of graphs to model a time-evolving scene. Finally, as in the previous study, interpretability of models and results is crucial to motivate further developments. A way to improve it is realizing a more robust analysis of the results by considering disaggregated results as done in Figure 15 and Figure 16. The distinction can be done on various criteria which add semantics, such as seasonality and rareness of the events, classification of the phenomenon or of the geographical site. Therefore, such an analysis should enhance the evaluation and comparison of different models and help explain the pros and cons of each model. We hope that this case study may help future works to better consider advantages and current limitations of graphs for a SITS forecasting task, thus encouraging the development of new methods.
@@ 6 Challenges and Perspectives
The abundance of data in Earth observation is the main cause of the scalability problem, with a constant flow of new data which are sparsely labelled. Various propositions of solution to handle this data volume exist, such as foundation models [233, 133] and continual learning. Exploring these solutions in the setting of spatio-temporal graphs could be promising. Adapting the spatio-temporal framework to dynamic scenarios to allow real-time updating, for example developing continuous-time dynamic graphs [99, 9] adapted to SITS, could also improve the scalability of the model.
@@
In this paper, graphs are principally used as tool to analyze and process SITS from a single source, mainly optical. Only a few of the aforementioned works have tackled other data sources, such as SAR [221, 42, 71], atmospheric [100, 109] and oceanic [119, 216] variables. However, Section 3.1 gave an overview of the wide range of sensors imaging the Earth. This diversity of sources brings development opportunities to adapt to each setting, even those with little data currently available (hyperspectral SITS, off-nadir and multi-angle acquisitions, 3D point clouds sequence). Graph-based approaches can also be adapted for data similar to SITS, such as aerial or drone imagery and numerical weather simulations.